A plane flying in a horizontal path 90m above the ground with a speed of 64 m/s², at what horizontal from the target should the pilot release the projectile?


1
Expert's answer
2021-01-14T10:39:53-0500

Let's first find the time that the projectile takes to reach the ground:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot90\ m}{9.8\ \dfrac{m}{s^2}}}=4.28\ s."

Finally, we can find the horizontal distance from the target at which the pilot should release the projectile:


"x=v_0t=64\ \dfrac{m}{s}\cdot4.28\ s=274\ m."

Answer:

"x=274\ m."